Circular No. 9075

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 S2 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery observa-

tion tabulated below) with a suggested westward tail ∼ 10′′ long on Sept. 20 UT on CCD images taken with the 0.5-m Uppsala Schmidt telescope; his images on Sept. 24.7 show the object to be slightly diffuse in good seeing with a several-arcsec extension to the west. After posting on the ‘NEOCP’ webpage, other CCD astrometrists have noted the object’s cometary appearance. E. Guido and G. Sostero (0.25-m reflector near Mayhill, NM, U.S.A., remotely; Sept. 24.4) write that co-added unfiltered exposures show a coma nearly 10′′ in diameter, elongated toward a narrow tail ∼ 14′′ long in p.a. 255^{o}. W. H. Ryan (Magdalena Ridge 2.4-m reflector; Sept. 24.45– 24.48) finds the object to be diffuse with a faint tail at p.a. ∼ 280^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Sept.20.67787 3^{h}33^{m}06.^{s}64 −24^{o}36′16′′.8 18.7 McNaught

The available astrometry (with Aug. 3 pre-discovery observations), elliptical orbital elements (T = 2009 June 23.6201 TT, e = 0.470488, q = 2.203599 AU, ω = 230^{o}.3660, Ω = 121^{o}.6146, i = 28^{o}.4493, equinox 2000.0; P = 8.49 years), and an ephemeris appear on MPEC 2009-S90.

1999 XY_{143}, 1999 RY_{214}, AND 2002 VT_{130}

K. S. Noll, Space Telescope Science Institute (STSI); W. M. Grundy, Lowell Observatory; S. D. Benecchi, STSI; and H. A. Levison, Southwest Research Institute, report the detection of three new transneptunian binaries. All of the observations were made with the Wide Field Planetary Camera 2 on the Hubble Space Telescope. In each instance, the pair of objects was detected in each of four dithered, 260-s exposures made through the F606W (wide-V bandpass) filter. The companion to 1999 XY_{143} (cf. MPEC 1999-Y19) was fainter by 0.38 magnitude, and was located −0′′.024 ± 0′′.001 in α and −0′′.082 ± 0′′.002 in δ, relative to the primary, in observations commencing on 2008 Sept. 19.1313 UT. The companion to 1999 RY_{214} (cf. MPECs 2000-A26, 2000-T17) was fainter by 1.09 magnitude, and was located +0′′.053 ± 0′′.006 in α and +0′′.017 ± 0′′.004 in δ, relative to the primary in observations commencing on 2008 Sept. 20.3703. The companion to 2002 VT_{130} (cf. MPECs 2002-X10 and 2004-B32) was fainter by 0.44 magnitude, and was located +0′′.042 ± 0′′.001 in α and +0′′.071 ± 0′′.002 in δ, relative to the primary in observations commencing on 2008 Sept. 21.9111.

2009 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green